OpenStudy (anonymous):

A curve in R^3 (set of real numbers) is generated by the intersection of the elliptic cylinder (x^2/2^2) + (y^2/3^2) = 1 and the PLANE z= (sq.rt(5)/2)x. a) Find the parametric representation for the curve with respect to a parameter t. b) Find the length of the curve from t=0 to t=pi/2.

OpenStudy (anonymous):

a) I'm not sure if this is the right way, but can't you say x = t and make your y(x) an y(x(t)) ? So this would mean that\[y(x(t)) = \sqrt{2^2+3^2-3^2t^2} / 2\]

OpenStudy (anonymous):

okay!! go on!

OpenStudy (amistre64):

that isnt parametricising the equation

OpenStudy (anonymous):

thats what i thought! i had to do something like.. x=cos t , y=sin t and something like that

OpenStudy (amistre64):

x = a cos(t) y = b sin(t) when x = 0; y = 3 which leads us to b=3 when y = 0; x = 2 which leads to a=2 x = 2 cos(t) y = 3 sin(t) z = (sqrt(5)/2) x z = (sqrt(5)/2) 2 cos(t) z = sqrt(5) cos(t)

OpenStudy (anonymous):

yes.. i got that far... but i am stuck there

OpenStudy (amistre64):

\[\int_{0}^{pi/2}\sqrt{[x'(t)]^2+[y'(t)]^2+[z'(t)]^2}\ dt\]

OpenStudy (anonymous):

yes... keep going please!

OpenStudy (amistre64):

trying to see if im missing any parts to it :) its been awhile since i tried these

OpenStudy (amistre64):

i am missing a part, just cant quite see how it fits in yet

OpenStudy (amistre64):

\[\int_{a}^{b}\ f\left(x(t),y(t),z(t)\right)\ \sqrt{[x'(t)]^2+[y'(t)]^2+[z'(t)]^2}\ dt\]

OpenStudy (amistre64):

its what to do with the f(.......) part i cant quite determine

OpenStudy (amistre64):

got it ...lol

OpenStudy (amistre64):

f(x,y,z) = \(\frac{x^2}{4} + \frac{y^2}{9} = 1\)

OpenStudy (amistre64):

well, its something to do with that

OpenStudy (anonymous):

oh.. okay...

OpenStudy (anonymous):

is there more to it or does it end there?

OpenStudy (amistre64):

there is more to it, but I cant focus at the moment, got my international business class soming up in a few

OpenStudy (anonymous):

oh.. okay... can you help me out when you can witht he same problem!

OpenStudy (anonymous):

just another question i am finding it difficult to figure out http://openstudy.com/groups/mathematics#/users/benfraser1012